OpenStudy (ambermarie151):

P(r)=r^2-0.1 144 What is the approximate value of P(3)? A.0.20 B.0.034 C.0.062 D.0.041

Directrix (directrix):

What is the 144 in the posted problem?

OpenStudy (ambermarie151):

It's supposed to be divide by 2-0.1/144

Directrix (directrix):

@ambermarie151 Would you click on the blue DRAW tab and write the equation as it appears in your text? Thanks.

jimthompson5910 (jim_thompson5910):

The equation is probably \[\Large P(r) = \frac{r^2-0.1}{144}\] but that's just an educated guess

OpenStudy (ambermarie151):

Yes that's it

jimthompson5910 (jim_thompson5910):

ok so you just replace every copy of r with 3 \[\Large P(r) = \frac{r^2-0.1}{144}\] \[\Large P({\color{red}{r}}) = \frac{({\color{red}{r}})^2-0.1}{144}\] \[\Large P({\color{red}{3}}) = \frac{({\color{red}{3}})^2-0.1}{144}\] then you evaluate the right side

jimthompson5910 (jim_thompson5910):

hopefully that makes sense @ambermarie151

OpenStudy (ambermarie151):

So would you do 3•3 -0.1 then divide by 144 ?

jimthompson5910 (jim_thompson5910):

correct since 3^2 means 3*3

OpenStudy (ambermarie151):

I am still confused

OpenStudy (ambermarie151):

I get a decimal

jimthompson5910 (jim_thompson5910):

what decimal value did you get?

jimthompson5910 (jim_thompson5910):

it's perfectly normal in this case to get a decimal value